A sum of $1350 is to be divided between two people in the ratio of 7 to 8. How much does each person receive?
Answer by mananth(16949)

$ 1350 is to be divided in the ratio of 7:8
It means one person gets 7 parts of $1350
and another gets 8 parts of $1350
7+8 = 15 parts make up $1350
7/15 is what the first person gets
7/15 * 1350
=7*90
=$630
8/15 * 1350= 90*8= 720
720+630= 1350